OpenStudy (loser66):

show that if n = 2phi(n) , where n is a positive integer, then n = 2^j for some positive integer j Please, help

OpenStudy (ikram002p):

same with previous question =) Hint:- n=2phi(n) phi(n)=n/2

OpenStudy (loser66):

Let \(n = p_1^{a_1}p_2^{a_2}...p_k^{a_k}\) \(\phi(n) = n \prod (\dfrac{p_i-1}{p_i})\) \(2\phi (n) = 2n\prod(\frac{p_i-1}{p_i})=n \implies 2\prod(\dfrac{p_i-1}{p_i}) =1\)

OpenStudy (loser66):

@ikram002p although they look alike but I don't think they want me to solve them the same. This is the problem next to the previous problem. If there is no different, why do they arrange them separately?

OpenStudy (loser66):

hence \(2(p_1-1)(p_2-1)...(p_k -1) = p_1p_2...p_k\) Then I stuck. ;)

OpenStudy (ikram002p):

i think it's only playing with minds :O . ok i'll figure other solution for u don't worry :)

ganeshie8 (ganeshie8):

Hint : \(p-1 \gt 1\) for all \(p\ne 2\)

OpenStudy (loser66):

then? how it forced p =2?

OpenStudy (ikram002p):

n is even as \( n=2\phi(n)\) , thus \(n = 2^{a_1}p_2^{a_2}...p_k^{a_k} \),all distinct primes. so for all numbers from 1 to n-1, we have half of them (even numbers) that are not relatively prime to n , so \(n>=2\phi(n)\), assume there exist a factor \(p_i \)of n is not 2 then there exist at least one odd number 1< j<n which is not relatively prime to n thus \(n>2\phi(n) \) which is a contradiction.

OpenStudy (loser66):

moreover, if \(p_i =2\) then the right hand side has the required form 2^j , but the left hand side is just 2.

ganeshie8 (ganeshie8):

that looks good to me

ganeshie8 (ganeshie8):

ikram's recent reply

OpenStudy (loser66):

Yes, I know, since it looks like the logic on the previous one which is phi ( n) = n/2, find n

OpenStudy (loser66):

I got it @ikram002p Thank you so much.

OpenStudy (ikram002p):

wait no i said i'm trying to figure out some other method -.- i posted this as remark

OpenStudy (ikram002p):

would it be right using contrapositive like, show that if \(n \neq 2^j\) for some integer j then \(n\neq 2\phi(n)\)

OpenStudy (ikram002p):

and u can use counterexample like \(n=p^j \) for odd p hmm idk u shouldn't close the question that fast -.- .

OpenStudy (loser66):

ok, let try one more time

OpenStudy (ikram002p):

ur earlier method \(2\phi (n) = 2n\prod(\frac{p_i-1}{p_i})=n \) so we need to show that \(2\prod(\dfrac{p_i-1}{p_i}) =1\)

OpenStudy (loser66):

n is even, hence \(n = 2^s p_1^{a_1}p_2^{a_2}...p_k^{a_k}\)

OpenStudy (loser66):

why? it is available there, just cancel n both sides, right?

OpenStudy (ikram002p):

consider primes set {2,3,5,7...} \(\dfrac {1}{2}\times \dfrac{2}{3}\times \dfrac {4}{5}....\times \frac{p_i-1}{p_i} \) is decreasing as p_i increase right ? so its like have a upper bound of 1/2 agree ?

OpenStudy (loser66):

hey, I don't get what you mean by saying that the sequence is decreasing . is it not that 1/2 < 2/3< 4/5<... how is it decreasing?

OpenStudy (ikram002p):

oh! i didn't say "sequence", i said the whole term \(\dfrac {1}{2}\times \dfrac{2}{3}\times \dfrac {4}{5}....\times \frac{p_i-1}{p_i} \) is decreasing.

OpenStudy (ikram002p):

nvm then its another stupid method, i should sleep now. goodnight.

OpenStudy (loser66):

Good night, friend. I do appreciate. :)

OpenStudy (loser66):

Oh, I gooooooot it hehehe \(2(p_1-1)(p_2-1)...(p_k -1) = p_1p_2...p_k\) but \(p_i -1\) is even, and \((p_i-1) | p_1p_2...p_k\) That forced each of p_i must be even. that is none of them is odd, hence n = 2^j

OpenStudy (ikram002p):

i logged in again to write that =) glad u got it.

OpenStudy (loser66):

:)